OpenStudy (anonymous):

please help!!!! Solve for x: cos(X) + cos(2x) = -1

OpenStudy (goformit100):

DO YOU KNOW THE FORMULA OF cos(2x) ????

OpenStudy (anonymous):

Here's a cheat sheet if you need it. http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

He means if you change cos(2x) to some function of cos(x), then you can more easily solve for cos(x).

OpenStudy (anonymous):

x=-1/2 cos^(-1)(-1-cos(X))

OpenStudy (anonymous):

and another solution is x=1/2 cos^(-1)(-1-cos(X))

OpenStudy (anonymous):

but what is x then??

OpenStudy (anonymous):

\[\cos(x)+\cos(2x)=-1\] \[\cos(x)+\cos^{2}(x)-\sin^{2}(x)+1=0\] \[\cos(x)+2\cos^{2}(x)=0 ...since...1-\sin^{2}(x)=\cos^{2}(x)\] \[\cos(x)=0...or...1+2\cos(x)=0\] Can U find x from here ?

OpenStudy (anonymous):

What interval are you solving over? There are four solutions between 0 and 2π.

OpenStudy (anonymous):

yes i can solve x from here...thanks you guys for helping!!